OpenStudy (anonymous):

I have a question, help me to solve this: A school sends 20 female members and 10 male members of the red society to provide first aid during the national day celebration. the number of male members who wear a hat is 6. if a member is chosen at random, the probability that he/she wears a hat is 3/10. calculate the number of female members who wear a hat.

OpenStudy (anonymous):

Figure out the total number of people who wear a hat.

OpenStudy (anonymous):

how can i figure that?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

The total number of people times the probability a person wears a hat.

OpenStudy (anonymous):

it's not 12

OpenStudy (anonymous):

How about this: \[ \frac{6+x}{10+20}=\frac{3}{10} \]

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

The total number of people who wear hats is \(6+x\) Where \(x\) is number of females with hat.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks wio